JOINT DIFFERENTIAL GEOMETRY/PDE AND PROBABILITY SEMINAR

Wednesday, February 16, 2005 Padelford C-36 3:50 pm Compact Heisenberg manifolds

Gerald Folland (UW)

Let M be the quotient of the (2n + 1)-dimensional Heisenberg group by a discrete co-compact subgroup; thus M is a strongly pseudoconvex compact CR manifold. We answer the following questions: (1) What are the eigenforms and eigenvalues of the Kohn Laplacians on M? (2) How can M be realized as the boundary of a bounded domain in a complex manifold? The answer to (1) plays a role in the answer to (2), which reveals an interesting connection with the theory of Abelian varieties.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.